### v2

# Emotic Masked Autoencoder with Attention Fusion for Facial Expression Recognition

Abstract
Facial Expression Recognition (FER) is a critical task within computer vision with diverse applications across various domains. Addressing the challenge of limited FER datasets, which hampers the generalization capability of expression recognition models, is imperative for enhancing performance. Our paper presents an innovative approach integrating the MAE-Face self-supervised learning (SSL) method and Fusion Attention mechanism for expression classification, particularly showcased in the 6th Affective Behavior Analysis in-the-wild (ABAW) competition. Additionally, we propose preprocessing techniques to emphasize essential facial features, thereby enhancing model performance on both training and validation sets, notably demonstrated on the Aff-wild2 dataset.
1 Introduction
Aff-wild2 [9, 8, 4, 3, 2, 13, 12, 7, 5, 11, 10, 6, 21] is a large-scale video database for ABAW competitions. It annotated 548 videos, around 2.7M frames, into eight pre-defined categories: anger, disgust, fear, happiness, sadness, surprise, neutral, and other. Thanks to the release of this database, we conduct experiments to explore the effectiveness of our method on this in-the-wild video database
Although facial expression recognition (FER) is a long-term and popular in computer vision tasks, it still have lots of limitations, such as lack of balance dataset, noise of synthetic faces affecting model performances, etc. So on, we determine to participate in solving these problems and our contribution can be mainly summarized as:
  • We develop a new technique for finetuning pretrained model MAE-Face to enhance model focus on more specific components of faces and emotions.
  • We promote new fusion model to integrate the latent features for boosting model ability and performances.
  • We experience our research into Expression Recognition task in ABAW 6th and get many promising results on validation set.
2 Related Work
Numerous innovative methods have emerged in previous editions of the Affective Behavior Analysis in-the-wild (ABAW) competitions. Our investigation delves into various expression classification and multi-task learning techniques, with a focus on the Expr task.
Zhang et al. [22] ensemble multiple 2D backbones to extract features for every single frame and concatenate these features to a temporal encoder to explore temporal features. By combining regression layers and classification layers, it learns from multi-task annotation and ranks first in the ABAW4 challenge. Li et al. [14] also harnessed MAE pre-trained weights in conjunction with AffectNet supervised pre-trained weights, securing a commendable 2nd place in ABAW4. Zhang et al. [23] introduced a transformer-based fusion module to integrate multi-modality features from audio, image, and textual data.
Ma et al. [16] achieved 1st place in the 5th ABAW challenge by employing MAE-Face alongside multiple-stage training. Jeong et al. [1] expanded the DAN model, resulting in a 2nd place finish in ABAW3. Xue et al. [19] implemented a coarse-to-fine cascade network with temporal smoothing, earning 3rd place in ABAW3.
3 Methodology
Figure 1: The architecture of our proposed model. The model consists of four components: pre-trained MAE-Face for feature extractor, fine-tuned MAE-face on Aff-wild2 for feature extractor, fusion blocks and MLP for final prediction.
In this section, we will describe our proposed approach in detail at Fig. 1. At the core of our model lies an architecture composed of several critical components. Initially, we employ a pre-trained Masked Autoencoder (MAE-Face) as a primary feature extractor to capture nuanced facial expressions and affective cues. This base is further refined by fine-tuning the MAE-Face on the Aff-wild2 dataset, ensuring our feature extraction is highly tailored to the intricacies of affective behavior. To process these features effectively, we incorporate self-attention blocks, which allow our model to discern the most relevant affective signals dynamically. The integration of these elements culminates in a Multilayer Perceptron (MLP) that serves as the final prediction mechanism, offering a precise and reliable analysis of emotional expressions.
3.1 Pre-processing
Figure 2: After the images are crop aligned, we extract the image parts which just contain mouth and eye for further training.
In the fine-tuning phases, we crop face images by percentage of frame sizes to get eye and mouth features. We take the eyes by cutting from 0.35 to 0.55 of height and 0.2 to 0.8 of width, we also crop the mouth with the same length but in 0.7 to 0.9 of height [Fig. 2]. Moreover, in order to uphold the integrity and quality of the pre-training dataset, we implement a rigorous data cleaning process. This involves the systematic removal of corrupt files, non-face images, and those with insufficient resolution, thereby ensuring the robustness and reliability of the dataset for subsequent analyses and model training.
3.2 MAE-Face
The proposed model employs a Vision Transformer (ViT) pre-trained via a self-supervised learning method, leveraging a Masked Autoencoder (MAE) for the process called MAE-Face[15]. Specifically, input images are segmented into 16$\times$16 patches, of which 75% are masked, focusing the training on reconstructing these patches from the remaining visible ones. The model is pre-trained on a comprehensive facial image dataset, including AffectNet [17], CASIA-WebFace[20], IMDB-WIKI[18], and CelebA[24], totaling 2,170,000 images, to enhance its capability in understanding facial features without relying on labels. Post pre-training, the model transitions to fine-tuning for downstream tasks Expression recognition on the Aff-wild2 dataset, then we extracted the feature base on the Pre-training and fine-tune MAE-face for further processing.
3.3 Fusion Attention block for Emotion Recognition
The following approach we employ a fusion-based approach for enhancing facial emotion recognition (FER) by integrating two pre-trained models. The methodology employs an attention-based network, focusing on both self-attention and local attention, to combine the strengths of these models effectively.
3.4 Fusion Attention Network
The proposed fusion attention network incorporates a multi-layer perceptron (MLP) for combining features from two emotion recognition models. By concatenating and then downsampling these features, the network applies self-attention and local attention mechanisms in the Multihead Attention Block to refine the feature representation for emotion classification.
3.4.1 Self-Attention and Local Attention
The self-attention mechanism operates by computing the attention scores based on the dot product of features from pre-trained networks, serving as the Key ($K$) and Query ($Q$), followed by a softmax function. The local attention mechanism employs convolution operations for extracting more refined features. The attention formula is defined as:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{K}}}\right)\otimes V$ (1)
where $K$ is the Key, $Q$ is the Query, $V$ is the Value, and $d_{K}$ is the dimensionality of the Key. This approach not only leverages the distinct advantages of both attention mechanisms but also enhances the model’s ability to recognize complex emotional states from facial expressions.
3.4.2 Loss function
Given the imbalance in the dataset for the facial expression classification challenge, we employ a weighted cross-entropy loss function. This decision is aligned with the need to handle the uneven distribution of the data across different classes. The loss function can be expressed as:
$\mathcal{L}_{\text{EXPR.CE}}=-\frac{1}{8}\sum_{j=1}^{8}W_{expr_{j}}\log\hat{y}_{j}$ (2)
By choosing this approach, we aim to give more focus to the minority classes, thereby improving the model’s performance on underrepresented data.
3.5 Post-processing
Given that the Aff-Wild2 dataset is derived from the entirety of video frames and facial expressions typically evolve gradually over time, rapid changes in expressions across adjacent frames are uncommon. Furthermore, following the removal of non-face, corrupted, and low-resolution images, a sliding window approach is adopted for post-processing prediction results to ensure smoother labels.
4 Experiment
In this section, we will provide a detailed description of the used datasets, the experiment setup, and the experimental results.
4.1 Dataset
For this ABAW challenge, Kollias et al.[21] collected a large-scale video database named Aff-Wild2. It consists of 548 videos and was labeled frame-by-frame. For the expression classification challenge, every frame in the video is annotated with one of eight pre-defined expression categories: anger, disgust, fear, happiness, sadness, surprise, neutral, and others. We randomly sampled 7000 images from each class for a small experiment dataset.
4.2 Setup
All training face images are resized to 224×224 pixels, our proposed method is implemented with the PyTorch toolbox on NVIDIA Tesla A6000 GPUs. The model is fine-tuned with the Adam optimizer for 100 iters. The learning rate for fine-tuning process is set to 1e-4. The batch size is set to 512. The average f1 Score across all eight categories on the validation set is reported.
Methods Results
MAE (Normal) 0.3596
MAE (Cropped) 0.3032
Fusion (Normal, Cropped) 0.3805
Table 1: Results on validation set
5 Conclusion
This paper employs a streamlined pipeline to assess the efficacy of the Fusion MAE-Face method for affective analysis on the Aff-Wild2 dataset. Our model enhances expression learning through improved preprocessing techniques, selectively isolating facial features to mitigate noise from surrounding pixels. Empirical investigations on Aff-Wild2 demonstrate that models initialized with MAE-Face pre-trained weights, coupled with fusion attention mechanisms, excel in extracting informative features, surpassing the performance of supervised approaches.